OpenStudy (anonymous):

Find the first and second derivatives of the function. y = xe^3x

OpenStudy (anonymous):

y'=3xe^3x + e^3x y'=(3x+1)e^3x y''=9xe^3x + 3e^3x + 3e^3x y''=9xe^3x + 6e^3x y''=3(3x+2)e^3x

OpenStudy (anonymous):

o i forgot to put the 3 infront of the ex^3x

OpenStudy (anonymous):

=)